A rock is thrown upward from level ground in such a way that the maximum height of its flight is equal to its horizontal range R. At what angle u is the rock thrown?
1
Expert's answer
2021-02-18T18:56:26-0500

considering the rock; "d=y_{max}"

"\frac{v^2sin2\theta}{g}=\frac{v^2sin^2\theta}{2g}"

"2sin\theta cos\theta=\frac{sin^2\theta}{2}"

"2cos\theta=\frac{sin\theta}{2}"

"tan\theta=4"

"\theta=tan^{-1}4=75.96^0"